OpenStudy (anonymous):

Graph the ellipse and locate the foci. 16x^2 +9y^2= 144

OpenStudy (luigi0210):

Do you know to start?

OpenStudy (anonymous):

is it x^2/a^2+ y^2/b^2=1

OpenStudy (luigi0210):

Yup, now do you know how to get this equation to equal one?

OpenStudy (anonymous):

16^x2/16+9y^2/16=144/16

OpenStudy (anonymous):

I am not sure

OpenStudy (luigi0210):

Hm, you sure about that dan?

OpenStudy (dan815):

i think so

OpenStudy (dan815):

i think i messed up y stuff

OpenStudy (anonymous):

do I need to use this c^2= a^2+b^2

OpenStudy (luigi0210):